OpenStudy (anonymous):

how do i get y by itself with 4x-4y=8

OpenStudy (anonymous):

x - y = 2 y = x - 2

hero (hero):

First divide each term by four, then add y to both sides while subtracting 2 from both sides to get: y = x - 2

OpenStudy (anonymous):

thankyou both !

hero (hero):

Just to warn you, the method will not always be the same. In this case, 4 just happens to be a common factor of itself and 8.